OpenStudy (anonymous):

If f(x)=-x^2-4x-k and contains the point (-4,3), what is the y coordinate of the function at the point on the axis of symmetry of the graph?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

These are NOT SAT practice problems.

Directrix (directrix):

f(x)=-x^2-4x-k and contains the point (-4,3), 3 = - (-4)^2 - 4(-4) - k @sakigirl Solve the above equation for k.

Directrix (directrix):

@sakigirl k = ?

Directrix (directrix):

>the point on the axis of symmetry The axis of symmetry is a line and is composed of infinitely points. So, I don't know what the reference to THE point is.

OpenStudy (anonymous):

the point is needed to calculate k

OpenStudy (anonymous):

f(x) = -x^2 - 4x - k at (-4, 3): 3 = -(-4)^2 -4(-4) - k 3 = -16 + 16 - k 3 = -k k = -3

OpenStudy (anonymous):

Now you'll have to convert this into a form where you know the vertex point(this will tell you what the axis of symmetry is).